# Hessian stability and convergence rates for entropic and Sinkhorn potentials via semiconcavity

### v1

@@ 1.3. Literature review
At the level of the primal solutions, namely entropic plans, let us mention [12] and [30]. In the former, the difference in (symmetric) entropy between the solutions to two different EOT problems is controlled in terms of a negative Sobolev norm, for a wide class of problems with costs induced by diffusions on Riemannian manifolds with Ricci curvature bounded from below (which includes the quadratic cost on $\mathbb{R}^{d}$). The latter obtains instead a quantitative Hölder estimate between the Wasserstein distance of optimal plans and that of their marginals. This result applies not only to the quadratic cost, but also to more general costs. We refer the reader to [12, Section 1.2.1] for a comparison between the two references. Let us further cite [34], where a more qualitative stability result is proven under mild hypotheses.
@@
Finally, [13] provides a control on the entropy between two entropic plans in terms of the (squared) Wasserstein distance between the marginals. The peculiarity of this last work is the approach, since it exploits for the first time the propagation of semiconcavity along HJB to obtain a quantitative stability result for primal solutions. The second-order quantitative stability bounds on entropic and Sinkhorn potentials that we will show in this manuscript build upon this previous contribution. For this reason and for sake of completeness, we prove the entropic stability estimate via semiconcavity also in the present manuscript, but we provide a different proof, based on the stochastic representation of Schrödinger bridges (see Theorem 2.1 below).
As concerns the dual solutions, i.e., entropic potentials, in [19] an $\mathrm{L}^{\infty}$-Lipschitz bound is obtained; it applies to multimarginal problems, but it requires either the space or the cost to be bounded. In [23] the $\mathrm{L}^{\infty}$-norm of the difference between entropic potentials associated to two EOT problems is controlled by the Wasserstein distance of order one between the corresponding marginals, using an approach based on Hilbert’s metric; but again, among their hypotheses, both the cost function and the marginals’ supports are assumed to be bounded. On the other hand, [14] succeeds in controlling the same difference with the Wasserstein distance of order two of the respective marginals, provided the cost is bounded with two bounded derivatives, i.e., $C^{2,\infty}$; if the regularity of the cost is higher, say $C^{k+2,\infty}$, then the $\mathrm{L}^{\infty}$-norm of the difference between entropic potentials can be replaced by the $C^{k,\infty}$-norm. Let us comment that the interest in higher-order stability results for entropic potentials is motivated by the fact that the gradient of entropic potentials provide good proxies for OT maps ([36, 43, 12] in unbounded settings and [48] in semidiscrete ones) and entropic estimates can be leveraged to obtain in the $T$ vanishing limit estimates for Kantorovich potentials and OT maps (e.g. [31, 21, 37]). In particular, in the very recent work [37] the authors rely on estimates for regularized potentials combined with gluing arguments in the vanishing $T$ limit, in order to get quantitative stability estimates for OT maps.
@@
Finally, in [27] the $\mathrm{L}^{2}$-norm of the difference of the gradients of entropic potentials is controlled in a Lipschitz way by the Wasserstein distance between the corresponding marginals by leveraging a functional inequality for tilt-stable probability measures, see [17] and [4, Lemma 3.21], and under the assumption that both entropic potentials have a bounded Hessian. The dependence of the Lipschitz constant on the regularization parameter is polynomial, thus improving on earlier results, and marginals may have unbounded support. Among those just mentioned, this is the closest contribution to ours, since the authors of [27] use Lipschitzianity of the Schrödinger maps (and hence concavity/convexity bounds for entropic potentials) in order to prove stability bounds for the gradient $\nabla\varphi^{\mu}$. Therefore, our work can be seen as an extension to second-order quantitative bounds. Moreover, the stability bounds for the gradients that we state here behave as theirs (our potentials and theirs differ from a multiplicative prefactor $-T$). In the compact setting we get the same asymptotic behavior in $R$ and $T$, whereas if we put ourselves in the Caffarelli’s setting (i.e., Hessian of marginals upper and lower bounded), then our general estimate would not depend on $T$ and would behave as their stability result when assuming bi-Lipschitzianity. We have not stated this corollary since already covered by their results and because assuming this extra regularity would not improve the dependence in $T$ in the Hessian stability estimate. Lastly, let us mention here that our approach can be extended also to weakly log-concave marginals, by building upon semiconcavity estimates obtained in [13].
@@
Sinkhorn’s algorithm. Contributions to Sinkhorn’s algorithm in the literature date back to the works of Sinkhorn [55] and Sinkhorn and Knopp [56], from which the algorithm takes its name. It was originally considered in a discrete setting framework for doubly stochastic matrices and the first exponential convergence result was given in [32, 8]. There, the authors studied the contraction properties of the algorithm in the Hilbert’s projective metric. After the seminal work of [22], which opened up to possible application of EOT to machine learning, multiple papers dealt with the convergence of the algorithm. Particularly in bounded settings (i.e., compact spaces or bounded costs) this has already been well established in [18, 26, 10]. In particular [18] obtained the first exponential convergence results in the continuous setting using the Hilbert’s metric approach. However, this approach provides rates that depend exponentially from the regularizing parameter $T$ and cannot be extended to unbounded settings.
@@
On the other hand, much less was known for unbounded settings (including the most iconic and simple quadratic cost setting with log-concave marginals). In fact, the only widely general known qualitative convergence result was due to Rüschendorf [49], who established qualitative convergence for Sinkhorn plans in relative entropy. Recently, this result has been improved in [47] where the authors have managed to show qualitative convergence on the primal and dual sides, under mild assumptions on cost and marginals. The first quantitative convergence result in unbounded settings we are aware of is [30], subsequently improved in [33], where the authors prove a polynomial convergence rate. These works are based on quantitative stability estimates for EOT and the insightful interpretation of Sinkhorn’s algorithm as a block-coordinate descent algorithm on the dual problem [39, 3, 38].
@@
Only very recently, it has been established the exponential convergence for unbounded costs and marginals. Up to the authors’ knowledge, the first contribution in this setting is given in [15], which studies the quadratic cost. There, the main result is that if the marginals are weakly log-concave and the regularization parameter $T$ is large enough, exponential convergence of the gradients of the iterates holds (and their results work for any $T>0$ for Gaussian marginals). This kind of convergence is particularly useful as $\nabla\varphi^{\mu}$ approximates the Brenier map in the $T\rightarrow 0$ limit, see [12, 48]. Moreover, this is the first contribution that has highlighted how geometric assumptions on the marginals (such as log-concavity) can be leveraged to improve the dependence in the convergence rates, from exponential to polynomial in $T$. Later, following similar considerations, [16] has improved the exponential convergence results in the bounded setting, showing that the exponential rate of convergence deteriorates polynomially in $T$. With regard to the unbounded setting, the article [28] has subsequently managed to construct a suitable version of Hilbert’s metric for general unbounded costs. In contrast with [15], exponential convergence is shown for all values of $T$, under a growth condition assumption. Roughly speaking, therein the author assumes that the tails of the marginals decay (strictly) faster than the cost function considered. When applied to the quadratic cost, this assumption does not completely cover log-concave distributions and their perturbations, leaving out Gaussian marginals for example.
@@
The first paper that finally has managed to provide exponential convergence rates in general (possibly unbounded) settings, working for any regularization parameter $T>0$ and with polynomial dependence in $T$, is [13]. There, together with our coauthors, we show how semiconvexity and semiconcavity bounds on Sinkhorn potentials can be leveraged to obtain exponential convergence. Our geometric approach is broadly general and covers as particular cases the bounded settings as well as the (anisotropic) quadratic costs, which includes, for instance, also the case when considering as cost function the transition kernel induced by an Ornstein–Uhlenbeck process (i.e., the framework of the Schrödinger bridge problem with a non-Gaussian reference process). The key observation employed there is that the semiconcavity of the function defined in (1.2) is enough to deduce quantitative stability estimates and exponential convergence rates depending on the semiconcavity parameter $\Lambda(\varphi^{\mu}_{0})$.
@@
Lastly, it is worth mentioning different contributions that over the past few years have focused on different asymptotic properties of Sinkhorn’s algorithm. Let us just mention [6] for the relation with Monge-Ampère equation, [24] for the construction of Wasserstein mirror gradient flows, [50] for construction of a Transformer variant inspired by Sinkhorn’s algorithm, and the very recent series of contributions [1, 2, 25] that focus on the relation of Schrödinger bridges and Sinkhorn’s algorithm with the Riccati matrix difference equations, and the impact of these results in the context of multivariate linear Gaussian models and statistical finite mixture models (including Gaussian-kernel density estimation). We conclude by mentioning the recent work [29], where the authors investigate the convergence of iterative proportional fitting procedures for a more general class of problems (which includes EOT), whose proof is based on strong convexity arguments for the dual problem, which particularly highlights the role of the geometric interplay between the subspaces defining the constraints.
@@
We would like to conclude this review by mentioning results that have inspired us or are related to ours. As we have already stated, our strategy is based on stochastic analysis and second-order estimates along Hamilton–Jacobi–Bellman equations. This approach has been initially introduced in [20] where Conforti has proved weak semiconcavity estimates for entropic potentials by studying how this property propagates along HJB equations. In [15] and [13], this has already been employed for proving the exponential convergence of Sinkhorn’s algorithm and for showing stability estimates of entropic plans. Here we further extend its use to show second-order stability estimates. In order to prove the convergence of Sinkhorn’s algorithm, a similar approach has been employed also in [35, 36] where we have studied how Lipschitzianity propagates along HJB equations, leading to a more perturbative convergence result (instead of a geometric one). Lastly, we would like to mention [11], though not directly applied to EOT, where the authors provide third-order estimates propagated along HJB in order to prove stability estimates for stochastic optimal control problems. These new ideas open up to further investigation of third-order estimates for the entropic potentials.
@@ 2. Preliminaries
In this paper we are interested in the behavior of the forward process $(Y^{\theta}_{s})_{s\in[0,T]}$ and backward process $(Y^{\eta}_{s})_{s\in[0,T]}$ defined as $Y^{\theta}_{s}\coloneqq\nabla\theta_{s}(X^{\psi^{\mu},\rho}_{s})$ where $\theta_{s}\coloneqq\psi^{\nu}_{s}-\psi^{\mu}_{s}$ and similarly as $Y^{\eta}\coloneqq\nabla\eta_{s}(X^{\varphi^{\mu},\mu}_{s})$ with $\eta_{s}\coloneqq\varphi^{\nu}_{s}-\varphi^{\mu}_{s}$. Since both $\varphi^{\nu}_{\cdot}$ and $\varphi^{\mu}_{\cdot}$ solve (HJB) it is immediate to see that $\eta_{\cdot}$ solves
$\partial_{s}\eta_{s}+\frac{1}{2}\Delta\eta_{s}-\nabla\varphi^{\mu}_{s}\cdot\nabla\eta_{s}-\frac{1}{2}|\nabla\eta_{s}|^{2}=0\,,$
and hence from Itô’s formula we may further deduce that
(2.1) $\mathrm{d}Y^{\eta}_{s}=\nabla^{2}\varphi^{\nu}_{s}(X^{\varphi^{\mu},\mu}_{s})Y^{\eta}_{s}\,\mathrm{d}s+\nabla^{2}\eta_{s}(X^{\varphi^{\mu},\mu}_{s})\,\mathrm{d}B_{s}\,.$
Similarly, when considering the backward processes we have
$\partial_{s}\theta_{s}+\frac{1}{2}\Delta\theta_{s}-\nabla\psi^{\mu}_{s}\cdot\nabla\theta_{s}-\frac{1}{2}|\nabla\theta_{s}|^{2}=0\,,$
and hence Itô’s formula implies
(2.2) $\mathrm{d}Y^{\theta}_{s}=\nabla^{2}\psi^{\nu}_{s}(X^{\psi^{\mu},\rho}_{s})Y^{\theta}_{s}\,\mathrm{d}s+\nabla^{2}\theta_{s}(X^{\psi^{\mu},\rho}_{s})\,\mathrm{d}B_{s}\,.$
Finally, notice that
$\mathbb{E}[|Y^{\eta}_{T}|^{2}]=\|\nabla\varphi^{\nu}-\nabla\varphi^{\mu}\|_{\mathrm{L}^{2}(\rho)}^{2}=\|\nabla\psi^{\nu}_{0}-\nabla\psi^{\mu}_{0}\|_{\mathrm{L}^{2}(\rho)}^{2}=\mathbb{E}[|Y^{\theta}_{0}|^{2}]\,.$
@@
Let us start by showing how the above stochastic control point of view can be employed in studying entropic stability for Schrödinger plans.
@@ 2.1. Entropic and gradients’ stability
Our Hessian stability result will rely on the following result regarding the stability of plans and gradients of (backward evolved) potentials.
@@
Assume H 1 and let $\pi^{\nu},\,\pi^{\mu}$ denote the optimal plans associated to EOT with marginals $(\rho,\nu)$ and $(\rho,\mu)$ respectively. Then it holds
$\mathscr{H}(\pi^{\mu}|\pi^{\nu})\leq\mathscr{H}(\mu|\nu)+\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,\mathbf{W}_{2}^{2}(\mu,\nu)\,.$
@@
(2.6) $\mathbb{E}_{\mu}[\mathscr{H}(\pi^{\mu}(\cdot|X)|\pi^{\nu}(\cdot|X))]\leq\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,\mathbf{W}_{2}^{2}(\mu,\nu)$
and
(2.7) $\mathbb{E}[|Y^{\eta}_{0}|^{2}]=\|\nabla\varphi^{\mu}_{0}-\nabla\varphi^{\nu}_{0}\|_{\mathrm{L}^{2}(\mu)}^{2}\leq\frac{\Lambda(\varphi^{\nu}_{0})\,C^{\varphi^{\nu}}}{T^{2}}\,\mathbf{W}_{2}^{2}(\mu,\nu)\,,$
@@
The entropic stability estimate stated above has already been proven by the authors and collaborators in [13, Theorem 1.1]. We report it here since the proof of this entropic stability estimate can be employed in order to get (2.6) and (2.7), which will play a crucial role in the rest of the paper. Finally, let us also remark that in here we provide a stochastic analysis proof of the entropic stability estimate by building a suitable competitor using a modified Schrödinger bridge process (see (2.10) below).
The above result will follow from combining the following technical bounds. In the first one we bound the relative entropy between $\pi^{\mu}_{s}(\cdot|y)=\mathcal{L}(X^{\varphi^{\mu},y}_{s})$ and $\pi^{\nu}_{s}(\cdot|z)=\mathcal{L}(X^{\varphi^{\nu},z}_{s})$.
[…]
@@
Next, since $(\nabla\varphi^{\nu}_{s}(X^{\varphi^{\nu},y}_{s}))_{s\in[0,T]}$ is a martingale (cf. [20, Proof of Theorem 2.1], namely it follows from Itô’s formula combined with (HJB) and (1.4)), we have
@@
so that if we integrate from 00 to $s$ the dynamics of $X_{s}^{\varphi^{\nu},y}$ and take expectations, we get
@@
with $g^{y}_{\varphi^{\nu}_{0}}$ defined as in (1.2), and we have used its $\Lambda(\varphi^{\nu}_{0})$-semiconcavity. ∎
In the particular case in which $s=T$ and we take $\mu=\nu$ (henceforth $\eta_{\cdot}=0$), the above result simply reads as
[…]
Assume H 1. For any $y,z\in\operatorname{supp}(\nu)$ we have
$\mathscr{H}(\pi^{\nu}(\cdot|y)|\pi^{\nu}(\cdot|z))\leq\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,|z-y|^{2}\,.$
The above corollary has already been proven by the authors and their collaborators in [13, Lemma 2.1], which can thus be seen as a particular case of the more general Section 2.1.
@@
Let us focus on the entropic stability bound first. Without loss of generality we may assume $\mathscr{H}(\mu|\nu),\,\Lambda(\varphi^{\nu}_{0}),\,\mathbf{W}_{2}^{2}(\mu,\nu)$ to be all finite, otherwise there is nothing to prove. Next, observe that $\pi^{\mu}$ can be seen as the entropic optimal plan w.r.t. the reference measure $\pi^{\nu}$ for the EOT problem

### v2

@@ 1.3. Literature review
At the level of entropic plans, let us mention [CCGT23] and [EN22]. In the former, the difference in (symmetric) entropy between the solutions to two different EOT problems is controlled in terms of a negative Sobolev norm, for a wide class of problems with costs induced by diffusions on Riemannian manifolds with Ricci curvature bounded from below (which includes the quadratic cost on $\mathbb{R}^{d}$). The latter obtains instead a quantitative Hölder estimate between the Wasserstein distance of optimal plans and that of their marginals. Let us further cite [GNB22], where a more qualitative stability result is proven under mild hypotheses. Finally, [CCGT24] provides a control on the entropy between two entropic plans in terms of the (squared) Wasserstein distance between the marginals. The peculiarity of this last work is the approach, since it exploits for the first time the propagation of semiconcavity along HJB to obtain a quantitative stability result for primal solutions. The second-order quantitative stability bounds on entropic and Sinkhorn potentials that we will show in this manuscript build upon this previous contribution. For this reason and for sake of completeness, we prove the entropic stability estimate via semiconcavity also in the present manuscript, but we provide a different proof, based on the stochastic representation of Schrödinger bridges (see Theorem 2.1 below).
@@
As concerns dual solutions, i.e., entropic potentials, in [CL20] an $\mathrm{L}^{\infty}$-Lipschitz bound is obtained; it applies to multimarginal OT, but it requires either the space or the cost to be bounded. In [DdBD24] the $\mathrm{L}^{\infty}$-norm of the difference between entropic potentials associated to two EOT problems is controlled by the Wasserstein distance of order one between the corresponding marginals, using an approach based on Hilbert’s metric; but again, both the cost function and the marginals’ supports are assumed to be bounded. On the other hand, [CCL24] succeed in controlling the same difference with the Wasserstein distance of order two of the respective marginals, provided the cost is $C^{2,\infty}$, i.e., bounded with two bounded derivatives; if the regularity of the cost is higher, say $C^{k+2,\infty}$, then the $\mathrm{L}^{\infty}$-norm of the difference between entropic potentials can be replaced by the $C^{k,\infty}$-norm. The interest in higher-order stability results for entropic potentials is motivated by the fact that their gradients provide good proxies for OT maps ([Gre24, MS25, CCGT23] in unbounded settings and [PNW21] in semidiscrete ones) and entropic estimates can be leveraged to obtain in the $T$ vanishing limit estimates for Kantorovich potentials and OT maps [FGP20, CP23, KLM25]. In particular, very recently [KLM25] rely on estimates for regularized potentials combined with gluing arguments in the vanishing $T$ limit, to get quantitative stability estimates for OT maps. Finally, in [DNWP25] the $\mathrm{L}^{2}$-norm of the difference of the gradients of entropic potentials is controlled in a Lipschitz way by the Wasserstein distance between the corresponding marginals, using a functional inequality for tilt-stable probability measures, see [CE22] and [BBD24, Lemma 3.21], and under the assumption that both entropic potentials have bounded Hessian. The dependence of the Lipschitz constant on the regularization parameter is polynomial, thus improving on earlier results, and marginals may have unbounded support. Among those just mentioned, this is the closest contribution to ours, since [DNWP25] use Lipschitzianity of the Schrödinger maps (and hence concavity/convexity bounds for entropic potentials) in order to prove stability bounds for $\nabla\varphi^{\mu}$. Therefore, our work can be seen as an extension to second-order quantitative bounds. Moreover, our stability bounds for the gradients behave as theirs (our potentials and theirs differ from a multiplicative prefactor T-T). In the compact setting we get the same asymptotic behavior in $R$ and $T$, whereas if we put ourselves in the Caffarelli’s setting (i.e., Hessian of marginals upper and lower bounded), then our general estimate would not depend on $T$ and would behave as their stability result when assuming bi-Lipschitzianity.
@@
Sinkhorn’s algorithm. Contributions to Sinkhorn’s algorithm in the literature date back to [Sin64] and [SK67]. It was originally considered in a discrete setting framework for doubly stochastic matrices and the first exponential convergence result was given in [FL89, BLN94] by exploiting properties of Hilbert’s projective metric. After the seminal work of [Cut13], which opened up to possible application of EOT to machine learning, multiple papers dealt with the convergence of the algorithm. Particularly in bounded settings (i.e., compact spaces or bounded costs) this has already been well established in [CGP16, DMG20, Car22]. In particular [CGP16] obtained the first exponential convergence results in the continuous setting using the Hilbert’s metric approach. However, this approach provides rates that depend exponentially on the regularizing parameter $T$ and cannot be extended to unbounded settings.
@@
On the other hand, much less was known for unbounded settings (including the most iconic and simple quadratic cost setting with log-concave marginals). In fact, the only widely general known qualitative convergence result was due to [Rus95], and it has been recently improved in [NW22]. The first quantitative result in unbounded settings we are aware of is [EN22], subsequently improved in [GN25], where the authors prove a polynomial convergence rate. These works are based on quantitative stability estimates for EOT and the insightful interpretation of Sinkhorn’s algorithm as a block-coordinate descent algorithm on the dual problem [Lég21, AFKL22, LAF23].
@@
Only very recently, it has been established the exponential convergence for unbounded costs and marginals. Up to the authors’ knowledge, the first contribution in this setting is given in [CDG23], which studies the quadratic cost. There, the main result is that if the marginals are weakly log-concave and the regularization parameter $T$ is large enough, exponential convergence of the gradients of the iterates holds (and their results work for any $T>0$ for Gaussian marginals). Moreover, this is the first contribution that has highlighted how geometric assumptions on the marginals (such as log-concavity) can be leveraged to improve the dependence in the convergence rates, from exponential to polynomial in $T$. Later, following similar considerations, [CDV25] improved the exponential convergence results in the bounded setting, showing that the exponential rate of convergence deteriorates polynomially in $T$. With regard to the unbounded setting, [Eck25] has subsequently managed to construct a suitable version of Hilbert’s metric for general unbounded costs. In contrast with [CDG23], exponential convergence is shown for all values of $T$, under a growth condition assumption. Roughly speaking, therein the author assumes that the tails of the marginals decay (strictly) faster than the cost function considered. When applied to the quadratic cost, this assumption does not completely cover log-concave distributions and their perturbations, leaving out Gaussian marginals for example.
@@
The first paper that has finally managed to provide exponential convergence rates in general (possibly unbounded) settings, working for any regularization parameter $T>0$ and with polynomial dependence in $T$, is [CCGT24]. There, together with our coauthors, we show how semiconcavity bounds on Sinkhorn potentials can be leveraged to obtain exponential convergence. Our geometric approach is broadly general and covers as particular cases the bounded settings as well as the (anisotropic) quadratic costs, which include, for instance, also the case when the cost function is the transition kernel induced by an Ornstein–Uhlenbeck process (i.e., the framework of the Schrödinger bridge problem with a non-Gaussian reference process). The key observation employed there is that the semiconcavity of the function defined in (1.3) is enough to deduce quantitative stability estimates and exponential convergence rates depending on the semiconcavity parameter $\Lambda(\varphi^{\mu}_{0})$.
@@
Lastly, it is worth mentioning different contributions that over the past few years have focused on different asymptotic properties of Sinkhorn’s algorithm. Let us just mention [Ber20] for the relation with Monge–Ampère equation, [DKPS23] for the construction of Wasserstein mirror gradient flows, [SABP22] for construction of a Transformer variant inspired by Sinkhorn’s algorithm, and the very recent series of contributions [ADMM24, ADMM25, DM25] that focus on the relation of Schrödinger bridges and Sinkhorn’s algorithm with the Riccati matrix difference equations, and the impact of these results in the context of multivariate linear Gaussian models and statistical finite mixture models. We conclude by mentioning the recent work [EL25], where the authors investigate the convergence of IPFP for a more general class of problems (which includes EOT), whose proof is based on strong convexity arguments for the dual problem, highlighting the role of the geometric interplay between the subspaces defining the constraints.
@@
We would like to conclude this review by mentioning results that have inspired us or are related to ours. As we have already stated, our strategy is based on stochastic analysis and second-order estimates along HJB equations. This approach has been initially introduced in [Con24], where Conforti has proved weak semiconcavity estimates for entropic potentials by studying how this property propagates along HJB equations. In [CDG23] and [CCGT24], this has already been employed for proving the exponential convergence of Sinkhorn’s algorithm and for showing stability estimates of entropic plans. Here we further extend its use to show second-order stability estimates. In order to prove the convergence of Sinkhorn’s algorithm, a similar approach has been employed also in [GNCD23, Gre24] where we have studied how Lipschitzianity propagates along HJB equations, leading to a more perturbative convergence result (instead of a geometric one). Lastly, we would like to mention [CC24], though not directly applied to EOT; there, the authors provide third-order estimates propagated along HJB in order to prove stability estimates for stochastic optimal control problems. These new ideas open up to further investigation of third-order estimates for entropic potentials.
@@ 2. Preliminaries
In this paper we are interested in the behavior of the forward process $(Y^{\theta}_{s})_{s\in[0,T]}$ and backward process $(Y^{\eta}_{s})_{s\in[0,T]}$ defined as $Y^{\theta}_{s}\coloneqq\nabla\theta_{s}(X^{\psi^{\mu},\rho}_{s})$ where $\theta_{s}\coloneqq\psi^{\nu}_{s}-\psi^{\mu}_{s}$ and similarly as $Y^{\eta}\coloneqq\nabla\eta_{s}(X^{\varphi^{\mu},\mu}_{s})$ with $\eta_{s}\coloneqq\varphi^{\nu}_{s}-\varphi^{\mu}_{s}$. Since both $\varphi^{\nu}_{\cdot}$ and $\varphi^{\mu}_{\cdot}$ solve (HJB) it is immediate to see that $\eta_{\cdot}$ and $\theta_{\cdot}$ solve
$\partial_{s}\eta_{s}+\frac{1}{2}\Delta\eta_{s}-\nabla\varphi^{\mu}_{s}\cdot\nabla\eta_{s}-\frac{1}{2}|\nabla\eta_{s}|^{2}=0\,,\qquad\partial_{s}\theta_{s}+\frac{1}{2}\Delta\theta_{s}-\nabla\psi^{\mu}_{s}\cdot\nabla\theta_{s}-\frac{1}{2}|\nabla\theta_{s}|^{2}=0\,.$
Hence from Itô’s formula we further deduce that
(2.1a) $\mathrm{d}Y^{\eta}_{s}=\nabla^{2}\varphi^{\nu}_{s}(X^{\varphi^{\mu},\mu}_{s})Y^{\eta}_{s}\,\mathrm{d}s+\nabla^{2}\eta_{s}(X^{\varphi^{\mu},\mu}_{s})\,\mathrm{d}B_{s}\,$,
(2.1b) $\mathrm{d}Y^{\theta}_{s}=\nabla^{2}\psi^{\nu}_{s}(X^{\psi^{\mu},\rho}_{s})Y^{\theta}_{s}\,\mathrm{d}s+\nabla^{2}\theta_{s}(X^{\psi^{\mu},\rho}_{s})\,\mathrm{d}B_{s}\,.$
Finally, notice that $\mathbb{E}[|Y^{\eta}_{T}|^{2}]=\|\nabla\varphi^{\nu}-\nabla\varphi^{\mu}\|_{\mathrm{L}^{2}(\rho)}^{2}=\|\nabla\psi^{\nu}_{0}-\nabla\psi^{\mu}_{0}\|_{\mathrm{L}^{2}(\rho)}^{2}=\mathbb{E}[|Y^{\theta}_{0}|^{2}]$.
@@ 2.1. Entropic and gradients’ stability
Let us start by showing how the stochastic control point of view can be employed in studying entropic stability of the optimal plans.
@@
Assume H 1 and let $\pi^{\nu},\,\pi^{\mu}$ denote the optimal plans associated to EOT with marginals $(\rho,\nu)$ and $(\rho,\mu)$ respectively. Then the following entropic stability bound holds
(2.5) $\mathscr{H}(\pi^{\mu}|\pi^{\nu})\leq\mathscr{H}(\mu|\nu)+\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,\mathbf{W}_{2}^{2}(\mu,\nu)\,.$
@@
(2.6a) $\mathbb{E}_{\mu}[\mathscr{H}(\pi^{\mu}(\cdot|X)|\pi^{\nu}(\cdot|X))]\leq\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,\mathbf{W}_{2}^{2}(\mu,\nu)\,$,
(2.6b) $\mathbb{E}[|Y^{\eta}_{0}|^{2}]=\|\nabla\varphi^{\mu}_{0}-\nabla\varphi^{\nu}_{0}\|_{\mathrm{L}^{2}(\mu)}^{2}\leq\frac{\Lambda(\varphi^{\nu}_{0})\,C^{\varphi^{\nu}}}{T^{2}}\,\mathbf{W}_{2}^{2}(\mu,\nu)\,$,
@@
The bound (2.5) has already been proven by the authors and collaborators in [CCGT24, Theorem 1.1]. We report it here since its proof can be employed in order to get (2.6a) and (2.6b), which will play a crucial role in the rest of the paper. Moreover, let us also remark that we provide here a stochastic analysis proof of (2.5) by building a suitable competitor using a modified Schrödinger bridge process (see (2.8) below). Before proving this result, let us state a technical bound, akin to [CCGT24, Lemma 2.1]. In particular the following lemma can be seen as a generalization of it, where we bound the relative entropy between $\pi^{\mu}_{s}(\cdot|y)=\mathcal{L}(X^{\varphi^{\mu},y}_{s})$ and $\pi^{\nu}_{s}(\cdot|z)=\mathcal{L}(X^{\varphi^{\nu},z}_{s})$, where we recall from (1.5) the backward Schrödinger bridge process (started in $x\in\mathbb{R}^{d}$) being defined by
(2.8) $\mathrm{d}X^{\varphi^{\nu},x}_{s}=-\nabla\varphi^{\nu}_{s}(X^{\varphi^{\nu},x}_{s})\mathrm{d}s+\mathrm{d}B_{s}\,,\quad X^{\varphi^{\nu},x}_{0}=x.$
Lemma 2.2.
@@
Next, since $(\nabla\varphi^{\nu}_{s}(X^{\varphi^{\nu},y}_{s}))_{s\in[0,T]}$ is a martingale (cf. [Con24, Proof of Theorem 2.1], namely it follows from Itô’s formula combined with the Hamilton–Jacobi–Bellman equation and the SDE of the Schrödinger bridge), we have
@@
so that if we integrate from 0 to $s$ the dynamics of $X_{s}^{\varphi^{\nu},y}$ and take expectations, we get
@@
with $g^{y}_{\varphi^{\nu}_{0}}(z)=\frac{|z-y|^{2}}{2}-T\,\varphi^{\nu}_{0}(z)$, and we have used its $\Lambda(\varphi^{\nu}_{0})$-semiconcavity. ∎
In the particular case $s=T$ and $\mu=\nu$ (henceforth $\eta_{\cdot}=0$), the above result simply reads as [CCGT24, Lemma 2.1], that is
Corollary 2.3.
Assume H 1. For any $y,z\in\operatorname{supp}(\nu)$ we have $\mathscr{H}(\pi^{\nu}(\cdot|y)|\pi^{\nu}(\cdot|z))\leq\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,|z-y|^{2}$.
@@
Let us focus on (2.5) first. Without loss of generality we may assume $\mathscr{H}(\mu|\nu),\,\Lambda(\varphi^{\nu}_{0}),\,\mathbf{W}_{2}^{2}(\mu,\nu)$ to be all finite, otherwise there is nothing to prove. Next, observe that $\pi^{\mu}$ can be seen as the entropic optimal plan w.r.t. the reference measure $\pi^{\nu}$ for the EOT problem